# 2024ICMLA.bib

### v1

@@ Abstract
This study explores the use of several LLMs for providing quantitative zero-shot sentiment analysis of implicit software desirability expressed by users. The study provides scaled numerical sentiment analysis unlike other methods that simply classify sentiment as positive, neutral, or negative. Numerical analysis provides deeper insights into the magnitude of sentiment, to drive better decisions regarding product desirability.
Data is collected through the use of the Microsoft Product Desirability Toolkit (PDT), a well-known qualitative user experience analysis tool. For initial exploration, the PDT metric was given to users of ZORQ, a gamification system used in undergraduate computer science education. The PDT data collected was fed through several LLMs (Claude Sonnet 3 and 3.5, GPT4, and GPT4o) and through a leading transfer learning technique, Twitter-Roberta-Base-Sentiment (TRBS), and through Vader, a leading sentiment analysis tool, for quantitative sentiment analysis. Each system was asked to evaluate the data in two ways, first by looking at the sentiment expressed in the PDT word/explanation pairs; and by looking at the sentiment expressed by the users in their grouped selection of five words and explanations, as a whole. Each LLM was also asked to provide its confidence (low, medium, high) in its sentiment score, along with an explanation of why it selected the sentiment value.
All LLMs tested were able to statistically detect user sentiment from the users’ grouped data, whereas TRBS and Vader were not. The confidence and explanation of confidence provided by the LLMs assisted in understanding the user sentiment. This study adds to a deeper understanding of evaluating user experiences, toward the goal of creating a universal tool that quantifies implicit sentiment expressed.
@@ I Introduction
    How do various LLMs compare to known sentiment analysis tools in providing sentiment analysis scores of PDT data?
@@
    Is there a particular algorithm (or combination of algorithms) that outperforms others and should be used as the basis for the development of an implicit user sentiment analysis tool?
@@ III-C Sentiment Analysis Tools and Methods
This research explores the sentiment analysis performance on PDT data of several LLMS: GPT4 [GPT4] and GPT4o [GPT4o], Claude Sonnet 3 and 3.5 [Claude], along with Twitter-Roberta-Base-Sentiment (TRBS) [Roberta], and Vader (Valence Aware Dictionary and sEntiment Reasoner) [Hutto_Gilbert_2014]. The latter two were chosen for their established effectiveness in sentiment analysis tasks and their differing approaches, which provide a useful comparison to the LLMs. TRBS is a pre-trained model fine-tuned on Twitter data for sentiment analysis, based on the RoBERTa architecture, and is a leading transfer learning technique. Vader, a long standing tool in this domain, combines a robust lexicon with heuristic rules for contextual nuance, making it user-friendly and highly accurate [36]. It is designed to perform well on social media text but effective across other text forms as well.
@@ V Discussion and Future Work
The results presented in Table II demonstrate that the LLMs performed well to quantify implicit user sentiment of PDT data using both respondent-level and Avg5 scoring. This research is an initial step toward the creation of a tool for a broad user-base designed to provide rich quantitative sentiment analysis of implicit product desirability, especially in situations where no user rating system exists. Based on the results, either GPT4o or Claude3.5 or a combination of both would serve well as the basis for this tool.

### v2

@@ Abstract
This study explores the use of LLMs for providing quantitative zero-shot sentiment analysis of implicit software desirability, addressing a critical challenge in product evaluation where traditional review scores, though convenient, fail to capture the richness of qualitative user feedback. Innovations include establishing a method that 1) works with qualitative user experience data without the need for explicit review scores, 2) focuses on implicit user satisfaction, and 3) provides scaled numerical sentiment analysis, offering a more nuanced understanding of user sentiment, instead of simply classifying sentiment as positive, neutral, or negative.
Data is collected using the Microsoft Product Desirability Toolkit (PDT), a well-known qualitative user experience analysis tool. For initial exploration, the PDT metric was given to users of two software systems. PDT data was fed through several LLMs (Claude Sonnet 3 and 3.5, GPT4, and GPT4o) and through a leading transfer learning technique, Twitter-Roberta-Base-Sentiment, and Vader, a leading sentiment analysis tool. Each system was asked to evaluate the data in two ways, by looking at the sentiment expressed in the PDT word/explanation pairs; and by looking at the sentiment expressed by the users in their grouped selection of five words and explanations, as a whole. Numerical analysis is used to provide insights into the magnitude of sentiment to drive high quality decisions regarding product desirability. Each LLM is asked to provide its confidence (low, medium, high) in its sentiment score, along with an explanation of its score.
All LLMs tested were able to statistically detect user sentiment from the users’ grouped data, whereas TRBS and Vader were not. The confidence and explanation of confidence provided by the LLMs assisted in understanding user sentiment. This study adds deeper understanding of evaluating user experiences, toward the goal of creating a universal tool that quantifies implicit sentiment.
@@ I Introduction
    How do various LLMs compare to known tools in providing sentiment analysis scores of PDT data?
@@
    Is there a particular algorithm that outperforms others and could be used as the basis for the development of an implicit user sentiment analysis tool?
@@ III-C Sentiment Analysis Tools and Methods
This research explores sentiment analysis performance on PDT data of several current LLMS: GPT4 [GPT4], GPT4o [GPT4o], Claude Sonnet 3 and 3.5 [Claude], along with Twitter-Roberta-Base-Sentiment (TRBS) [Roberta], and Vader (Valence Aware Dictionary and sEntiment Reasoner) [Hutto_Gilbert_2014]. The latter two were chosen for their established effectiveness in sentiment analysis tasks and their differing approaches, which provide a useful comparison to the modern LLMs. TRBS is a pre-trained bidirectional model fine-tuned on Twitter data for sentiment analysis, based on the RoBERTa architecture, and is a leading transfer learning technique. Vader, a long standing tool in this domain, combines a robust lexicon with heuristic rules for contextual nuance, making it user-friendly and highly accurate. It was designed to perform well on social media text but is effective across other text forms. The selected LLMs (GPT4o and Claude 3.5) represent two of the most advanced LLMs [Krugmann], and our model is adaptable for use with other LLMs.
@@ V Discussion and Future Work
The results presented in Table II demonstrate that the LLMs performed well to quantify implicit user sentiment of PDT data using both respondent-level and Avg5 scoring. Additionally, similar results were found on the CARMA dataset. This research is an initial step toward the creation of a tool for a broad user-base designed to provide rich quantitative sentiment analysis of implicit product desirability, especially in situations where no user rating system exists. Based on the results, either GPT4o or Claude3.5 or a combination of both would serve well as the basis for this tool. A combination could work similar to weighted voting or other bagging-like [bagging] machine learning methods.